OpenStudy (anonymous):

please HELP ...using implicit differentiation to find dy/dx given xy+cos(y)=x

OpenStudy (dumbcow):

just use normal derivative rules but remember whenever you take derivative of term with "y" you must multiply by dy/dx

OpenStudy (dumbcow):

\[\rightarrow y + x \frac{dy}{dx}-\sin(y) \frac{dy}{dx} = 1\] solve for dy/dx